Please help me solve the equation 3 sqrt(50)/2 sqrt(27)
Found 2 solutions by Fombitz, stanbon:
Answer by Fombitz(32388)
Answer by stanbon(75887)

solve the equation
3 sqrt(50)/2 = sqrt(27)
----------
3*5*sqrt(2)/2 = 3*sqrt(3)
----------
5sqrt(2)/2 = sqrt(3)
===================
Comment:: Your "equation" has no variable to solve for.
-------------------
Cheers,
Stan H.
-------------------